Any help would be greatly appreciated


Solve x^2 - 7x - 3 = 0 for x using any method you prefer
Thanks
Holly

Answer by richwmiller(17219)

x^2 - 7x - 3 = 0
12/4=3
(7/2)^2+3=49/4+12/4
complete the square (x-7/2)^2=61/4
I leave the rest for you. You should have some of the fun.